# Push and Pull: A Framework for Measuring Attentional Agency

### v1

@@ 5.1. Advertising in the Technology Industry
A common refrain in early criticisms of the targeted advertising model was “if you are not paying for the product, you are the product.” Subsequent proposals, such as data as labor (Posner and Weyl, 2018) and data dignity (Lanier and Weyl, 2018), pursue a related critique: that under the current paradigm, users of free internet platforms are also effectively workers producing data. Advocates of the advertising model, on the other hand, argue that it promotes equity and access because free products extend the benefits of technological innovation to those at the bottom of the socio-economic ladder.
Our framework shows how key claims frequently made in these debates can be made concrete. Our measures quantify the amount of attentional agency these platforms accord to various groups of individuals. If someone is not paying for a costly service, their activity is being subsidized in some way, often by advocates. The degree to which this reduces their attentional agency, however, depends on the relative alignment of preferences, technological context, and other factors.
This creates an open challenge for balancing the welfare implications of attentional agency and the consumer surplus afforded by ad-funded internet platforms (Brynjolfsson et al., 2023). Below, we outline policy ideas aimed at making tangible progress on this challenge, especially in light of recent AI progress.
@@ 5.2. Attentional Agency and Alignment
Our measures of attentional agency also complement ongoing research into AI alignment, especially work that emphasizes the challenge of aligning systems to the expressed or revealed values of diverse groups of stakeholders, using frames such as pluralism in human values (Sorensen et al., 2024) and social choice (Conitzer et al., 2024). Platforms create value by extending attention to previously unconsidered informational objects, which necessarily entails entrusting them with a certain degree of control over one’s attention. For AI-powered search and other systems that primarily function in this capacity, pull measures alignment to an individual user’s intentions. Push, on the other hand, measures alignment to other objectives, such as public trust and safety.
However, both push and pull represent special cases of more general conceptions of agency that are beginning to arise as AI systems become progressively more capable of handling decisions beyond information retrieval. A primary benefit of autonomous systems—i.e., ones that can act without needing to be continuously monitored—is that they free up people’s attention for other uses. This, however, necessarily reduces the amount of oversight such systems receive, which means they require substantially more trust in the faithful execution of user intentions and observance of public trust and safety objectives.
The concepts of push and pull clarify a variety of issues related to alignment and safety. For instance, interventions that aim to block specific concepts or pieces of information (e.g., instructions to build a weapon), such as preference ranking optimization (Song et al., 2024), or chained “safety classifiers” (Kim et al., 2023), and more can be understood as seeking to fulfill the goals of an advocate—which might be a public body acting in the public interest. In other words, alignment and safety interventions that are not highly personalized are best understood as supporting pushes from an advocate, often a representation of the public’s expectation of trust and safety. This insight is useful because modeling such situations as conflicts between advocates allows us to bring economic reasoning and mechanism design to bear on the problem.
@@ 5.3. Policy Strategies
There have been many proposals to promote greater transparency and accountability online. Some of the most prominent examples include the transparency provisions in the EU’s Digital Services Act (DSA), EU’s General Data Protection Regulation (GDPR), and proposed EU Artificial Intelligence Act (Economic and Committee, 2021; gdp, 2016)  [^6] Reporting of metrics that we outline in this paper are not new. For instance, Article 26 of the DSA requires platforms to provide users with meaningful information about the main parameters used in targeted advertising and ways to modify these settings. Article 26 also calls for a ban on targeted advertising based on profiling using special categories of personal data identified in Article 9 of the GDPR. Additionally, Article 27 of the DSA requires platforms to clearly explain the main parameters that influence what information is suggested to users and provide options to modify these parameters. Furthermore, Article 28 requires the largest online platforms and search engines to provide users with at least one recommender system option that is not based on profiling of individuals by automated systems as defined in the GDPR (Tech Policy Press, 2022).
Below, we highlight three policy strategies naturally suggested by our framework. Although a full implementation would require additional effort to account for platform-specific interactions with the framework, each of these relatively simple proposals stand to improve the distribution of informational agency in a measurable way.
@@
    Transparency into attentional agency: Measure and provide information on the distributions of Pull and Push values, both in aggregate and by subgroups. These disclosures would reveal particularly acute instances of misalignment between advocate and agent value on platforms. Greater transparency would be especially helpful for ensure that identifiable populations (e.g., people of a particular sexual orientation, racial identity, or age category) are experiencing different levels of attentional agency on a platform.
@@
    Expanded consumer choice: A variety of platforms sell push in order to generate revenue. For those that do, users could be given the option to pay for a “push-free” experience (which would generalize the “ad-free” tier offered by some services). An additional benefit of such options is that the revealed preferences of users who pursue it could inform standards for acceptable $\lambda$ values.
@@
While paid push-free options, neutral algorithm defaults, or push/pull disclosures may seem like relatively modest interventions, they reflect a fundamentally different theory of platforms than conventional transparency approaches. Rather than focusing on the black-box mechanics of ranking and recommendation systems, they target the concrete attentional choices that platforms make on behalf of users and the structural incentives that shape those choices. Disclosure-focused policy would open the door to broader societal conversations about equitable attentional agency.
These ideas are not a substitute for existing algorithmic audits, privacy protections, content moderation standards, or other key planks of the online safety agenda. Attentional agency attacks the problem of manipulative platform design from a neglected angle, one grounded in a numerical accounting of how platform incentives distort the information that users encounter. As policymakers grapple with the expanding domains in which algorithmic systems shape individual and collective experience, the push/pull framework helps identify and measure the core concept of attentional agency.
5.4. Impacts of Generative Foundation Models
Throughout this paper, we have commented on how our framework clarifies the likely effects of models that enable platforms to better understand users and to parse more granular units of information from across an increasing variety of sources. Given that these developments stand to disrupt the existing economic arrangement of the internet, it is also important to highlight how the above policy recommendations will interact with the present trajectory of AI development.
First, it could be the case that businesses avoid incorporating generative models into many of the platforms described above, in which our above recommendations remain highly relevant. Another possibility is that the ad-based funding model reasserts itself; perhaps ad-based search engines successfully integrate generative models while relying primarily on ad-based revenue, not subscriptions. In this scenario, transparency into $\lambda$ could become even more important, as ads become simultaneously more persuasive and less obviously ads. Finally, it could be the case that the rise of generative models catalyzes a general switch to subscription-based access to information. Of note, many leading AI efforts have chosen to rely on subscription and direct purchasing of API access. In this case, firms could see a stronger incentive to offer subscriptions, as this would be reflected in their informational agency disclosures.
Revenue generated by users paying to consume information could then be used to fund content creators. Although there are concerns in the short-term about potentially calamitous effects of language models on industries like journalism (Grynbaum and Mac, 2023), there exists a hopeful future in which direct access to highly personalized and carefully ranked chunks of information is worth paying for and ultimately leads to more funding allocated to information creation and curation. In this case, content creators may also become less reliant themselves on advertising. Achieving this hopeful future is, of course, highly path dependent, and will be impacted by regulatory outcomes in the near future and the ability for various information-providing industries to bargain with platforms that mediate information flow.
@@ 5.5. The Political Economy of Social Media and AI
Recent policy debates have highlighted a variety of issues that arise when large organizations have the capacity to determine what people pay attention to at scale. As one example, concerns about the political independence of TikTok’s content and algorithmic recommendations spurred the United States to pass legislation banning any “website, desktop application, mobile application, or augmented or immersive technology application” that is “controlled by a foreign adversary” (tik, 2024).
Our framework informs to these debates by providing a quantitative tool that can be used to evaluate the impact that competing incentives of various stakeholders have on user agency. By urging platforms to disclose push and pull metrics, policymakers can assess the actual balance of attentional agency on the platform and make evidence-based decisions about appropriate responses to emerging concerns.
The TikTok ban is just one example of the growing desire for better evidence-based approaches to platform accountability and governance. As just discussed in Section 5.4, generative foundation models and other innovations are quickly expanding the capacities of platforms, raising the stakes of developing a more participatory role for users in technological development (Allen and Weyl, 2024). By demanding transparency into the attentional agency of users, policymakers and the public can work towards aligning platform incentives with democratic values and the genuine preferences of users.
@@ 6. Conclusion
The framework we propose provides a lens for reasoning about how current platforms distribute attention. Our methods also provide new insight into ways that future platforms, especially ones powered by highly capable forms of artificial intelligence, might come closer to achieving the original visions of thinkers who inspired the internet. This framework we propose is intended to be readily applicable to policy discussions—indeed, the operational definitions we propose are based on techniques and computations already in use by platform operators.

### v2

@@ 5.1. Advertising in the Technology Industry
A common refrain in early commentary on the targeted digital advertising model was “if you are not paying for the product, you are the product.” Subsequent proposals, such as data as labor (Posner and Weyl, 2018) and data dignity (Lanier and Weyl, 2018), pursue a related critique: that under the current paradigm, users of free internet platforms are also effectively workers producing data. Advocates of the advertising model, on the other hand, argue that it promotes equity and access because free products extend the benefits of technological innovation to those at the bottom of the socio-economic ladder.
Our framework shows how key claims frequently made in these debates can be made more concrete by quantifying the amount of attentional agency these platforms accord to various groups of individuals. If someone is not paying for a costly service, their activity is being subsidized in some way, often by advocates. The degree to which this reduces their attentional agency, however, depends on the relative alignment of preferences, technological context, and other factors. Measures of push and pull show how the combined influence of these factors can be operationalized using relatively lightweight calculations.
This creates an open challenge for balancing the welfare implications of attentional agency and the consumer surplus afforded by ad-funded internet platforms (Brynjolfsson et al., 2023). Below, we outline ideas aimed at making tangible progress on this challenge, especially in light of recent AI progress.
@@ 5.2. Attentional Agency and Alignment
Our measures of attentional agency also complement ongoing research into AI alignment, especially work that applies concepts such as pluralism in human values (Sorensen et al., 2024) and social choice (Conitzer et al., 2024). Platforms create value by extending a user’s attention to unseen informational objects, which necessarily entails the user entrusting the platform with a certain degree of discretion. Thus, pull can measure alignment to an individual’s intentions. Push, on the other hand, measures alignment to the objectives of others.
Push and pull further clarify interventions that aim to block specific concepts or pieces of information (e.g., instructions to build a weapon), such as preference ranking optimization (Song et al., 2024) and chained “safety classifiers” (Kim et al., 2023). In our framework, such measures can be understood as seeking to fulfill the goals of an advocate—which might be a public body acting in the public interest. This insight is useful because modeling such situations as conflicts between advocates allows us to bring economic reasoning and mechanism design to bear on the problem.
5.3. Attentional Agency and Generative Foundation Models
Our framework reveals that, from the perspective of attentional agency, the key innovation of generative foundation models lies in their ability to parse information from a wide range of sources at an exceptionally granular level—that of the individual token. As we argued in Section 3.2.3, this enables AI-powered search tools to extract information from websites while leaving behind embedded advertising and other elements that support monetization. This raises a critical question: will some form of the advertising model endure, or will it give way to a new paradigm in which agents pay directly (e.g., via subscriptions or API credits) for access to information?
The potential for generative foundation models to circumvent existing monetization pathways has raised concerns about their implications for journalism and related industries (Grynbaum and Mac, 2023). Our framework highlights, however, that this problem arises precisely because these technologies enable agents to maximize pull. However, there exists a hopeful future in which direct access to highly personalized and carefully ranked chunks of information is worth paying for and ultimately leads to more funding for information creation and curation.
In short, if we view advances in AI as reducing the granularity of informational objects that can be processed on digital platforms, the lens of attentional agency gives us a clear vocabulary (and with regulatory support, concrete data—see below for how lightweight regulatory interventions could allow for the sharing and use of push and pull measurements) to reason about when the reduction in granularity will work to advance human values, and when it will create new challenges (e.g., violate previous arrangements around incentives).
@@ 5.4. Policy Strategies
There have been many proposals to promote greater transparency and accountability online. Some of the most prominent examples include the transparency provisions in the EU’s Digital Services Act (DSA), EU’s General Data Protection Regulation (GDPR), and proposed EU Artificial Intelligence Act (European Economic and Social Committee, 2021; EU Parliament, 2016)  [^12] Reporting of metrics that we outline in this paper are not new. For instance, Article 26 of the DSA requires platforms to provide users with meaningful information about the main parameters used in targeted advertising and ways to modify these settings. Article 26 also calls for a ban on targeted advertising based on profiling using special categories of personal data identified in Article 9 of the GDPR. Additionally, Article 27 of the DSA requires platforms to clearly explain the main parameters that influence what information is suggested to users and provide options to modify these parameters. Furthermore, Article 28 requires the largest online platforms and search engines to provide users with at least one recommender system option that is not based on profiling of individuals by automated systems as defined in the GDPR (Tech Policy Press, 2022).
Below, we highlight three policy strategies that naturally complement our framework. Notably, each of these simple proposals can be implemented with relatively little effort or tailoring by platforms.
@@
    Transparency into attentional agency: Measure and provide information on the distribution of Pull and Push, both in aggregate and by subgroup. Greater transparency would help users identify instances where they are experiencing reduced levels of attentional agency and make informed choices between platforms, potentially increasing both accountability and competition.
@@
    Expanded consumer choice: A variety of platforms sell push in order to generate revenue. For those that do, users could be given the option to pay for a “push-free” experience (which would generalize the “ad-free” tier offered by some services). This would enable users who place a greater premium on pull to maximize it for a price, increasing overall economic efficiency.
@@
These strategies all target the concrete attentional choices that platforms make on behalf of users and the structural incentives that shape those choices. These ideas are not a substitute for existing algorithmic audits, privacy protections, content moderation standards, or other key planks of the online safety agenda. Attentional agency attacks the problem of manipulative platform design from a neglected angle, one grounded in a numerical accounting of how platform incentives distort the information that users encounter.
While paid push-free options, neutral algorithm defaults, or push/pull disclosures may seem like relatively modest interventions, they reflect a fundamentally different theory of platforms than conventional transparency approaches. Rather than focusing on the black-box mechanics of ranking and recommendation systems, they target the concrete attentional choices that platforms make on behalf of users and the structural incentives that shape those choices. Importantly, measures to increase transparency at this level would support broader societal conversations about how attentional agency is distributed without requiring detailed technical knowledge of how algorithms produce specific rankings.
As we pointed out in our discussion of the “Economic, Psychological, and Social Value of Attention” (Section 2.4), people’s revealed attentional preferences are not always tightly coupled to their underlying welfare. Policies that uncritically seek to maximize pull may therefore not only fail to improve, but in fact actively undermine, a user’s overall health and happiness (e.g., if they are struggling with attentional self-control problems). One promising solution would be to develop and apply our framework to algorithms that infer people’s reflective (rather than impulsive) preferences (see (Kleinberg et al., 2024) for a general discussion).
@@ 5.5. The Political Economy of Social Media and AI
Recent policy debates have highlighted a variety of issues that arise when large organizations have the capacity to determine what people pay attention to at scale. As one example, concerns about the political independence of TikTok’s content and algorithmic recommendations spurred the United States to pass legislation banning any “website, desktop application, mobile application, or augmented or immersive technology application” that is “controlled by a foreign adversary” (U.S. Congress, 2024).
Our framework helps inform these debates by providing a quantitative tool that can be used to evaluate the impact that competing incentives of various stakeholders have on user agency. As just discussed in Section 5.3, generative foundation models and other innovations are quickly expanding the capacities of platforms, raising the stakes of developing a more participatory role for users in technological development (Allen and Weyl, 2024).
These developments reflect a growing desire for better evidence-based approaches to platform accountability and governance. By demanding transparency into the attentional agency of users, policymakers and the public can work towards aligning platform incentives with democratic values and the genuine preferences of users.
Measures of attentional agency also complement recent work on the interplay of algorithmic platforms and phenomena such as ideological segregation (González-Bailón et al., 2023), echo chambers and “rabbit holes” (Brown et al., 2022), and interaction with misinformation more generally (Edelson et al., 2021). This body of work provides a complementary lens for understanding the health of an algorithmic platform or feed. Given that push and pull measurements can be assessed in a faceted manner, it could be useful to consider push and pull along axes such as political valence or through the lens of “trust.”
@@ 6. Conclusion
The framework we propose provides a lens for reasoning about how current platforms distribute attention. Our methods also provide new insight into ways that future platforms, especially ones powered by highly capable forms of artificial intelligence, might come closer to achieving the original visions of the thinkers who inspired the internet. The framework we propose is intended to be readily applicable to policy discussions—indeed, the operational definitions we propose are based on techniques and computations already in use by platform operators.